find an equation of the line that contains the given pair of points.
(0,7) and (7,0)

Answer by stanbon(75887)

y-intercept is 7 since (0,7) is on the line.
slope = (0-7)/(7-0) = -1
EQUATION:
y=-x+7
Cheers,
Stan H.